# Small area estimation of growing stock timber volume, basal area, mean stem diameter, and stem density for mountain forests in Austria

### v1

@@ 1 Introduction
Planning and management occur at the stand-level; hence, they need to be informed by stand-level structure parameter estimates. Given cost constraints, installing sufficiently dense forest inventory plot networks to achieve the desired level of estimate precision using traditional design-derived estimators is typically not feasible. As a result, various approaches have been developed to estimate parameters in stands with insufficient numbers of observations. Such approaches commonly couple sparsely sampled forest inventory and remote sensing data within a non-parametric or parametric modeling framework. The efficiency of non-parametric nearest neighbor methods (e.g., Moeur and Stage,, 1995; Tomppo and Halme,, 2004; LeMay and Temesgen,, 2005; Maltamo et al.,, 2006), various interpretations of small area estimators (e.g., Goerndt et al.,, 2011, 2012; Breidenbach and Astrup,, 2012; Mandallaz et al.,, 2013; Steinmann et al.,, 2013; Mandallaz,, 2014; Green et al.,, 2019; Katila and Heikkinen,, 2020; Coulston et al.,, 2021; Green et al.,, 2022), and geostatistical methods (e.g., Mandallaz,, 2000; Tuominen et al.,, 2003; Zhang et al.,, 2004; Berterretche et al.,, 2005; Buddenbaum et al.,, 2005; Maselli and Chiesi,, 2006; Babcock et al.,, 2018; Finley et al.,, 2024) are explored in many studies and remain an active area of research.
@@
In our current study, extreme sparseness of the available forest inventory plots precludes stand-level direct estimates and hence use of area-level models. Rather, we pursue unit-level SAE using Bayesian regression models. Predictive performance is assessed for series of candidate models of increasing complexity. Data used to inform estimates are described in Section 2.1 and 2.2, the model definitions, their implementation, and prediction algorithms are given in Section 2.3 and 2.4. Candidate models and assessment metrics are described in Section 2.5. Results are given in Section 3, followed by a discussion and next steps are in Section 4.
@@ 2.2 Outcome and predictor variables
$\text{N}=\frac{\text{BA}}{\frac{\pi}{4}\cdot\left(\frac{\text{QMD}}{100}\right)^{2}}$ (4)
@@ 2.4.3 Posterior inference
A total of 500 MCMC batches with 10 iterations each were generated. The first 50% of the resulting 10,000 samples were discarded as burn-in, and the remaining samples were thinned by a factor of 10, yielding $T=250$ nearly independent draws.
@@
For each stand, the posterior mean and coefficient of variation (CV, expressed as a percentage) were computed for GSV, BA, QMD, and N. These summaries were spatially mapped, and the empirical cumulative distribution functions of CV were plotted to evaluate the distribution of relative prediction uncertainty across stands.
@@ 2.5.1 Evaluating candidate models
To assess model properties—specifically bias, accuracy, and precision—we employ cross-validation. A naive approach would use random unit-level cross-validation. While informative for evaluating unit-level predictive performance, this approach risks overestimating the performance of spatial models, as predictions for a held-out unit may be spatially close to observed units, leading to information leakage and inflated accuracy metrics.
To address this, spatial blocking is used to define cross-validation folds. In this approach, units are grouped into spatially coherent blocks, and all units within a block are held out simultaneously. This ensures that predictions are made for spatially distinct areas, aligning more closely with the objective of stand-level inference. By withholding entire spatial blocks, the influence of spatial autocorrelation on model performance metrics is reduced.
Spatial blocks were constructed using the blockCV R package (Valavi et al.,, 2019), with block sizes set to approximately 250 meters—the average distance between stand centroids and the nearest plot. This scale aligns the folds with the spatial resolution of the target inference. A 20-fold cross-validation was implemented, where each fold comprised a disjoint set of spatial blocks that contained approximately 7 plots. Models were fit using the remaining blocks, and predictions were generated for the held-out units. These predictions were aggregated to the block level following the procedure in Section 2.4.3, treating folds as stand analogues.
@@
Cross-validation performance was summarized using several metrics: bias, root mean squared prediction error (RMSPE), mean standardized bias and RMSPE (expressed as percentages), empirical coverage of 95% credible intervals (i.e., the proportion of held-out values captured by their corresponding 95% PPD intervals), and the average width of these intervals.
@@ 3.3 Cross-validation
While not our primary inferential objective, it may be instructive to consider the spatially blocked cross-validation metrics computed for unit-level predictions, which are presented in Table A.2. These scores are useful for evaluating models intended for point prediction within stands. The unit-level metric scores exhibit patterns consistent with those observed in the aggregated block-level results shown in Table 5. In particular, bias is generally low across models; the inclusion of predictor variables improves both accuracy and precision; and the addition of spatial random effects—whether univariate or multivariate—yields only marginal, if any, performance gains. At the unit level, where predictions directly align with the model structure, we would expect approximately 95% nominal coverage from the 95% credible intervals, which is indeed observed across all models.
@@ 3.4 Spatial prediction
To further assess predictive uncertainty, empirical cumulative distribution functions (ECDFs) of the stand-level CVs were computed separately for each outcome (Figure 5). The CV was less than 25% for 79.61% of GSV predictions, all QMD predictions, 84.1% of BA predictions, and 15.78% of N predictions.
Figure 5: Empirical cumulative distribution functions of the coefficient of variation for each outcome variable, computed across forest stands in Brixen im Thale.
4 Discussion and conclusion
@@
A key finding is that the spatial dependence observed in our study exhibited relatively short effective ranges, typically around or below 1 km. Much of the spatial structure in the raw measurement data is already accounted for by the ALS-derived predictors, leaving little residual spatial autocorrelation for the random effects to capture. Consequently, the spatial models yield only modest improvements over non-spatial counterparts, as the opportunity for spatial information borrowing is limited.

### v2

@@ 1 Introduction
Planning and management occur at the stand-level; hence, they need to be informed by stand-level structure parameter estimates. Given cost constraints, installing sufficiently dense forest inventory plot networks to achieve the desired level of estimate precision using traditional design-derived estimators is typically not feasible. As a result, various approaches have been developed to estimate parameters in stands with insufficient numbers of observations. Such approaches commonly couple sparsely sampled forest inventory and remote sensing data within a non-parametric or parametric modeling framework. The efficiency of non-parametric nearest neighbor methods (e.g., Moeur and Stage,, 1995; Tomppo and Halme,, 2004; LeMay and Temesgen,, 2005; Maltamo et al.,, 2006), various interpretations of small area estimators (e.g., Goerndt et al.,, 2011, 2012; Breidenbach and Astrup,, 2012; Mandallaz et al.,, 2013; Steinmann et al.,, 2013; Mandallaz,, 2014; Mauro et al.,, 2016; Breidenbach et al.,, 2018; Green et al.,, 2019; Katila and Heikkinen,, 2020; Coulston et al.,, 2021; Green et al.,, 2022), and geostatistical methods (e.g., Mandallaz,, 2000; Tuominen et al.,, 2003; Zhang et al.,, 2004; Berterretche et al.,, 2005; Buddenbaum et al.,, 2005; Maselli and Chiesi,, 2006; Babcock et al.,, 2018; Finley et al.,, 2024) are explored in many studies and remain an active area of research.
@@
In our current study, extreme sparseness of the available forest inventory plots precludes stand-level direct estimates and hence use of area-level models. Rather, we pursue unit-level SAE using Bayesian regression models. Predictive performance is assessed for series of candidate models of increasing complexity. Data used to inform estimates are described in Section 2.1 and 2.2, the model definitions, their implementation, and prediction algorithms are given in Section 2.3 and 2.4. Candidate models and assessment metrics are described in Section 2.5. Results are given in Section 3, followed by the discussion and conclusion in Sections 4 and 5, respectively.
@@ 2.1 Study region and data
Based on the maximum tree height and walking path in the area, we estimate the distance between the scanner and the tops of the tree crowns to be about 40 m. The PLS device has an initial beam size of 12.7 mm horizontally and 9.5 mm vertically, with beam divergences of 1.2 mrad and 3 mrad, respectively (Hyper Tech Advanced Technologies Ltd.,, 2017). These dimensions result in an elliptical footprint of approximately 0.95 cm2 at 0 m and 45.2 cm2 at 40 m. While the larger footprint in the upper parts of the tree reduces the ability to resolve small structures, the uncertainty caused by beam divergence ($\pm$6 cm at 40 m) does not significantly affect the tree height measurement. In fact, the increased footprint makes it more likely that the laser will hit small branches, thereby preserving the completeness of the point cloud, especially in the upper sections of the crowns. However, the accuracy of tree height measurements depends heavily on stand structure. In denser, multilayered forests, occlusion becomes a significant issue as laser pulses are hindered from reaching the treetops. Nonetheless, this problem is far less pronounced for PLS than for TLS, as PLS captures data continuously along the path, unlike TLS, which is limited to a few fixed scan positions (Compeán-Aguirre et al.,, 2024). In temperate forests, the accuracy of PLS-derived tree heights is comparable to traditional measurement methods (Jurjević et al.,, 2020). However, these results likely do not apply to tropical rainforests.
@@ 2.2 Outcome and predictor variables
$\text{N}=\frac{\text{BA}}{\frac{\pi}{4}\left(\frac{\text{QMD}}{100}\right)^{2}}$ (4)
@@ 2.4.3 Posterior inference
A total of 500 MCMC batches with 10 iterations each were generated. The first 50% of the resulting 5,000 samples were discarded as burn-in, and the remaining samples were thinned by a factor of 10, yielding $T=250$ nearly independent draws.
@@
For each stand, the posterior mean and coefficient of variation (CV, expressed as a percentage) were computed for GSV, BA, QMD, and N and mapped.
@@ 2.5.1 Evaluating candidate models
To assess model properties—specifically bias, accuracy, and precision—we employ cross-validation. A naive approach would use random plot-level cross-validation. While informative for evaluating unit-level predictive performance, this approach risks overestimating the performance of spatial models, as predictions for a held-out plots may be spatially close to observed plots, leading to information leakage and inflated accuracy metrics.
To address this, spatial blocking is used to define cross-validation folds. In this approach, plots are grouped into spatially coherent blocks, and all plots within a block are held out simultaneously. This ensures that predictions are made for spatially distinct areas, aligning more closely with the objective of stand-level inference. By withholding entire spatial blocks, the influence of spatial autocorrelation on model performance metrics is reduced.
Spatial blocks were constructed using the blockCV R package (Valavi et al.,, 2019), with block sizes set to approximately 250 meters—the average distance between stand centroids and the nearest plot. This scale aligns the folds with the spatial resolution of the target inference. A 20-fold (where folds are blocks) cross-validation was implemented, where each fold comprised a disjoint set of spatial blocks that contained approximately 7 plots. Models were fit using the remaining blocks, and predictions were generated for the held-out plots. These predictions were aggregated to the block level following the procedure in Section 2.4.3, treating blocks as stand analogues.
@@
Cross-validation performance was summarized using
$\text{bias}=\frac{1}{L}\sum_{l=1}^{L}\left(y_{q}^{\ast}(\mathcal{B}_{l})-y_{q}(\mathcal{B}_{l})\right)$ (10)
and root mean squared prediction error
$\text{RMSPE}=\sqrt{\frac{1}{L}\sum_{l=1}^{L}\left(y_{q}^{\ast}(\mathcal{B}_{l})-y_{q}(\mathcal{B}_{l})\right)^{2}},$ (11)
[…]
@@ 3.3 Cross-validation
While not our primary inferential objective, it may be instructive to consider the spatially blocked cross-validation metrics computed for unit-level predictions, which are presented in Table A.2. These scores are useful for evaluating models intended for point prediction within stands and to assess the appropriateness of the posited models for these unit-support data. The unit-level metric scores exhibit patterns consistent with those observed in the aggregated block-level results shown in Table 5. In particular, bias is generally low across models; the inclusion of predictor variables improves both accuracy and precision; and the addition of spatial random effects—whether univariate or multivariate—yields only marginal, if any, performance gains. At the unit level, where predictions directly align with the model structure, we would expect approximately 95% nominal coverage from the 95% credible intervals, which is indeed observed across all models.
These unit-level cross-validation metrics also provide insight into the appropriateness of the posited models. In particular, when defining the spatial processes, we assumed stationary GPs. If such assumptions about the error structure were grossly inappropriate, we would expect to see systematic deviations from nominal predictive coverage. However, the observed coverage rates remain consistently close to 95% across all models, suggesting the spatial variance components are well estimated and the stationarity assumption is not leading to biased or overconfident uncertainty estimates. While this does not preclude the possibility of localized departures from stationarity, these results suggest the stationary GP assumption is sufficiently flexible and performs well for the purposes of unit-level prediction and uncertainty quantification in this context.
Formal tests for non-stationarity do exist (see, e.g., Fuentes,, 2005), but they typically require large spatial domains and dense sampling to reliably detect departures from stationarity—conditions that are not met in the present setting. Even if moderate departures from stationarity were present, it is unclear whether introducing a more complex error structure would yield meaningful gains in predictive performance.
@@ 3.4 Spatial prediction
4 Discussion
@@
A key finding is that the spatial dependence observed in our study exhibited relatively short effective ranges, typically around or below 1 km. Much of the spatial structure in the raw measurement data is already accounted for by the ALS-derived predictors, leaving little residual spatial autocorrelation for the random effects to capture. Consequently, the spatial models yield only modest improvements over non-spatial counterparts, as the opportunity for spatial information borrowing is limited. Related studies have reported similar findings regarding the spatial structure of residuals after accounting for remotely sensed predictors. For example, Mauro et al., (2017) evaluated spatial autocorrelation in regression residuals when predicting forest volume, biomass, and QMD using ALS-derived covariates in central Spain. They found that these predictors explained a substantial portion of the spatial structure, leaving only short-range residual autocorrelation. Likewise, Breidenbach et al., (2016) fit spatial regression models to Norwegian National Forest Inventory data using photogrammetrically derived canopy height as a predictor for stand volume. After accounting for stand structure covariates, they observed very short residual spatial ranges—typically less than 25 meters. These findings reinforce the idea that much of the large-scale spatial dependence in forest inventory outcomes can be captured through appropriate predictor variables, leaving only localized residual structure to be modeled stochastically.